# Top-Bottom Interference Contribution to Fully-Inclusive Higgs Production

### v1

@@ Abstract
We evaluate the top-bottom interference contribution to the fully-inclusive Higgs production cross section at next-to-next-to-leading order in QCD. Although bottom-quark-mass effects are power-suppressed, the accuracy of state-of-the-art theory predictions makes an exact determination of this effect indispensable. The total effect of the interference is pb, while the pure correction is 0.43 pb. With this result, we address one of the leading theory uncertainties of the cross section.
@@ I Introduction
where the numerical value was calculated using GeV and GeV. Still, missing finite-quark-mass effects at NNLO remain a source of uncertainty. In particular, two contributions were each estimated to amount to in Ref. de Florian et al. (2016): the exact dependence on the top mass in all diagrams where the top quark remains the only quark which couples to the Higgs, and interference effects between such top-quark diagrams and diagrams containing a bottom or charm loop. These finite-quark-mass effects are generally power-suppressed: In the case of finite top-mass effects, the cross section contributions are suppressed by powers of close to the production threshold. For bottom quarks, on the other hand, they are suppressed by independently of partonic center-of-mass energy. The latter suppression originates partly from the Yukawa coupling constant and partly from the scalar nature of the interaction vertex, which requires a helicity flip of the quark. Helicity conservation then demands that the cross section vanishes as , resulting in an overall suppression of . However, the cross section also receives strong logarithmic enhancements of the form , rendering the total interference quite sizeable. Charm-quark contributions receive an analogous suppression, which is however even stronger due to the smaller mass value. Top- and bottom-quark-mass effects are hence viewed to be among the leading theory uncertainties. Recently, the production cross section was computed with full top-mass dependence at NNLO Czakon et al. (2021), confirming the expected size of the top-mass effects. So far, predictions of the fully inclusive top-bottom interference cross section are still lacking. Threshold expansions around at NLO and next-to-next-to-leading logarithmic resummations in the HTL Anastasiou and Penin (2020) estimated the NNLO contribution to the interference to be in the range of pb to 0.08 pb. On top of that, differential cross-section analyses for Higgs+jet production at NLO Grazzini and Sargsyan (2013); Melnikov and Penin (2016); Lindert et al. (2017); Caola et al. (2018); Bonciani et al. (2023) have become available and confirmed the significance of bottom quark effects, especially in the low- region. In these studies, the transverse momentum of the Higgs serves as an infrared regulator. In the case of the fully inclusive cross section, the possibility of a vanishing Higgs transverse momentum has to be taken into account. This necessitates the inclusion of double-virtual amplitudes and also a numerically stable implementation of real-radiation amplitudes in infrared-sensitive regions of the phase space.
@@ II Calculation
All contributions have been included in Stipper, the C++ implementation of the sector-improved residue subtraction scheme Czakon (2010); Czakon and Heymes (2014); Czakon et al. (2019), in order to correctly deal with all their infrared divergences. The cancellation of infrared divergences is checked by verifying that all poles in the dimensional regulator are compatible with zero within their Monte Carlo uncertainty.
The masses of the heavy quarks are renormalized in the on-shell scheme, while the strong coupling constant is renormalized in the -scheme with five light quark flavors. The Lehman-Symanzik-Zimmermann (LSZ) constant for the gluon field is non-trivial already at the one-loop order Czakon et al. (2008). The LSZ constants for the massless quark fields are non-trivial starting at two-loops, which makes them irrelevant for the present calculation. Since the bottom quark is assumed to be massless whenever the fermion loop is not coupled to the Higgs boson, it does not contribute to the gluon LSZ constant, and no decoupling constant needs to be introduced.
@@ III Results
Order | [pb] | [pb] | [pb] | [%] |
TeV | ||||
LO | – | |||
+9.14 | ||||
NLO | ||||
+4.19 | +0.167(3) | |||
NNLO | ||||
TeV | ||||
LO | – | |||
+21.14 | ||||
NLO | ||||
+9.72 | +0.147(1) | +0.434(8) | ||
@@
Table 1 presents the key findings of this publication. For completeness, we reproduce the HEFT results and the effects of a finite top-quark mass, whose value is taken as GeV, as in Ref. Czakon et al. (2021). Additionally, we provide scale uncertainties, which, in the case of the NNLO top-quark mass effects, constitute an additional novel result of this work. The main result is the top-bottom interference effect at the same top-quark mass and a bottom-quark mass of GeV, also including scale uncertainties. A calculation at a lower value, GeV, yields a compatible with the result given in Table 1 within the Monte Carlo uncertainty. For at 13 TeV, the contribution changes slightly from pb to pb, which is compensated by a shift at , leading to a negligible change of the NNLO result from pb to pb. Thus, we conclude that the effect of small variations of the top-quark mass value is too small to be sensibly quantified. While the size of the NNLO correction to is as estimated in Refs. de Florian et al. (2016); Anastasiou et al. (2016), our result exhibits some surprising properties. First, one notices that the top-bottom interference has an even larger effect than the exact top-mass dependence compared to HEFT. This is, however, not too surprising because the HEFT values already account for some of the top-mass effects with the rescaling procedure. Second, the corrections have opposite sign and similar magnitude compared to the contribution, leading to an almost complete cancellation. Furthermore, they are more than twice as large as expected from the NLO scale uncertainties. Together with the fact that we observe an increased scale dependence at NNLO compared to NLO, we can conclude that for top-bottom interference effects in Higgs production, standard scale variation at NLO underestimates the magnitude of higher-order contributions. Finally, the contribution to of pb can be compared to Ref. Anastasiou and Penin (2020), where the authors use a next-to-leading logarithmic approximation to estimate this effect to lie between pb and pb, implying that large effects are unaccounted for when this approximation is used.
@@ IV Conclusions and Outlook
We have computed the top-bottom interference contribution to the Higgs production cross section in gluon-gluon fusion at NNLO. The result addresses one of the most important remaining theory uncertainties for this process. We find that the effect is sizable and, therefore, crucial for precision predictions at the percent level. The result is compatible with conservative estimates previously available in the literature. Interestingly, the scale uncertainties grow slightly from NLO to NNLO. It has been noticed Anastasiou et al. (2016); Caola et al. (2018) that the interference can be very sensitive to the choice of the renormalization scheme. A dedicated analysis comparing the calculation presented in this letter to one in the scheme or the 4-flavor scheme would be desirable to analyze if there is a setup which shows a particularly good convergence behavior. We defer this question to future work. Due to these open questions regarding the scheme choice, we refrain from providing values for top-charm interference, even though the methods used in this work can be readily applied to obtain them. It seems sensible to first address the impact of scheme choice before considering this very small effect, where the usage of the 5-flavor scheme as explained above might no longer be appropriate.

### v2

@@ Abstract
We evaluate the top-bottom interference contribution to the fully-inclusive Higgs production cross section at next-to-next-to-leading order in QCD. Although bottom-quark-mass effects are power-suppressed, the accuracy of state-of-the-art theory predictions makes an exact determination of this effect indispensable. The total effect of the interference at 13 TeV is pb, while the pure correction is 0.43 pb. With this result, we address one of the leading theory uncertainties of the cross section.
@@ I Introduction
where the numerical value was calculated using GeV and GeV. Still, missing finite-quark-mass effects at NNLO remain a source of uncertainty. In particular, two contributions were each estimated to amount to in Ref. de Florian et al. (2016): the exact dependence on the top mass in all diagrams where the top quark remains the only quark which couples to the Higgs, and interference effects between such top-quark diagrams and diagrams containing a bottom or charm loop. These finite-quark-mass effects are generally power-suppressed: In the case of finite top-mass effects, the cross section contributions are suppressed by powers of close to the production threshold. For bottom quarks, on the other hand, they are suppressed by independently of partonic center-of-mass energy. The latter suppression originates partly from the Yukawa coupling constant and partly from the scalar nature of the interaction vertex, which requires a helicity flip of the quark. Helicity conservation then demands that the cross section vanishes as , resulting in an overall suppression of . However, the cross section also receives strong logarithmic enhancements of the form , rendering the total interference quite sizeable. Charm-quark contributions receive an analogous suppression, which is however even stronger due to the smaller mass value. Top- and bottom-quark-mass effects are hence viewed to be among the leading theory uncertainties. Recently, the production cross section was computed with full top-mass dependence at NNLO Czakon et al. (2021), confirming the expected size of the top-mass effects. So far, predictions of the fully inclusive top-bottom interference cross section are still lacking. A threshold expansion around at NLO with next-to-next-to-leading logarithmic resummation in the HTL Anastasiou and Penin (2020) has been used to estimate the interference to contribute -2.18 pb at NNLO. On top of that, differential cross-section analyses for Higgs+jet production at NLO Grazzini and Sargsyan (2013); Melnikov and Penin (2016); Lindert et al. (2017); Caola et al. (2018); Bonciani et al. (2023) have become available and confirmed the significance of bottom-quark effects, especially in the low- region. In these studies, the transverse momentum of the Higgs serves as an infrared regulator. In the case of the fully inclusive cross section, the possibility of a vanishing Higgs transverse momentum has to be taken into account. This necessitates the inclusion of double-virtual amplitudes and also a numerically stable implementation of real-radiation amplitudes in infrared-sensitive regions of the phase space.
@@ II Calculation
All contributions have been included in Stripper, the C++ implementation of the sector-improved residue subtraction scheme Czakon (2010); Czakon and Heymes (2014); Czakon et al. (2019), in order to correctly deal with all their infrared divergences. The cancellation of infrared divergences is checked by verifying that all poles in the dimensional regulator are compatible with zero within their Monte Carlo uncertainty.
The masses of the heavy quarks are renormalized in the on-shell scheme, while the strong coupling constant is renormalized in the scheme with five light quark flavors. The Lehmann-Symanzik-Zimmermann (LSZ) constant for the gluon field is non-trivial already at the one-loop order Czakon et al. (2008). The LSZ constants for the massless quark fields are non-trivial starting at two-loops, which makes them irrelevant for the present calculation. Since the bottom quark is assumed to be massless whenever the fermion loop is not coupled to the Higgs boson, it does not contribute to the gluon LSZ constant, and no decoupling constant needs to be introduced.
@@ III Results
Order | [pb] | [pb] | [pb] | [pb] | [%] |
TeV | |||||
LO | – | ||||
NLO | |||||
NNLO | |||||
TeV | |||||
LO | – | ||||
NLO | |||||
NNLO | |||||
TeV | |||||
LO | – | ||||
NLO | |||||
NNLO | |||||
TeV | |||||
LO | – | ||||
NLO | |||||
NNLO | |||||
TeV | |||||
LO | – | ||||
NLO | |||||
@@
Table 1 presents the key findings of this publication. For completeness, we reproduce the HEFT results and the effects of a finite top-quark mass, whose value is taken as GeV, as in Ref. Czakon et al. (2021). Additionally, we provide scale uncertainties, which, in the case of the NNLO top-quark mass effects, constitute an additional novel result of this work. The main result is the top-bottom interference effect at the same top-quark mass and a bottom-quark mass of GeV, also including scale uncertainties. A calculation at a lower value, GeV, yields a compatible with the result given in Table 1 within the Monte Carlo uncertainty. For at 13 TeV, the contribution changes slightly from pb to pb, which is compensated by a shift at , leading to a very small change of the NNLO result from pb to pb. Thus, we conclude that the effect of small variations of the top-quark mass value is too small to be sensibly quantified. While the size of the NNLO correction to is as estimated in Refs. de Florian et al. (2016); Anastasiou et al. (2016), our result exhibits some surprising properties. First, one notices that the top-bottom interference has an even larger effect than the exact top-mass dependence compared to HEFT. This is, however, not too surprising because the HEFT values already account for some of the top-mass effects with the rescaling procedure. Second, the corrections have opposite sign and similar magnitude compared to the contribution, leading to an almost complete cancellation. Furthermore, they are more than twice as large as expected from the NLO scale uncertainties. Together with the fact that we observe an increased scale dependence at NNLO compared to NLO, we can conclude that for top-bottom interference effects in Higgs production, standard scale variation at NLO underestimates the magnitude of higher-order contributions.
Finally, our result for at 13 TeV can be compared to Ref. Anastasiou and Penin (2020), where the authors calculate it, albeit at a very large top-quark mass value using the scheme for the bottom-quark mass and Yukawa coupling, using a next-to-leading logarithmic approximation and obtain pb at LO, pb at NLO, and pb at NNLO. Since their NNLO value lies almost within our scale uncertainty, we can confirm the quality of their approximation. Additionally, we observe a compelling reduction of the renormalization-scheme dependence.
To further investigate convergence and scheme dependence, we transformed our results to a mixed renormalization scheme, where the bottom mass stays on shell, while the Yukawa coupling is renormalized in the scheme. Contrary to a complete calculation with both the mass and the Yukawa coupling in the scheme, the mixed scheme is straightforward to implement since . The Yukawa-coupling running was obtained with CRunDec Schmidt and Steinhauser (2012) and the required scheme-translation constants can be found in Ref. Gray et al. (1990). The results are shown in Table 1. At NNLO, they are compatible with the on-shell results within scale uncertainties, albeit with a much weaker scale dependence. The convergence across different orders of perturbation theory is substantially improved, indicating a potentially better convergence of a consistent -scheme result. At present, our on-shell results with their larger errors seem appropriate as a conservative estimate for phenomenological applications.
@@ IV Conclusions and Outlook
We have computed the top-bottom interference contribution to the Higgs production cross section in gluon-gluon fusion at NNLO. The result addresses one of the most important remaining theory uncertainties for this process. We find that the effect is sizable and, therefore, crucial for precision predictions at the percent level. The result is compatible with conservative estimates previously available in the literature. Interestingly, the scale uncertainties grow slightly from NLO to NNLO. It has been noticed Anastasiou et al. (2016); Caola et al. (2018) that the interference can be very sensitive to the choice of the renormalization scheme. When changing only the renormalization of the bottom Yukawa coupling to , compatible results with weaker scale dependence are obtained. Since it is questionable if such a mixed renormalization scheme is consistent in the context of the complete Standard Model, and since the standard in the Higgs Working Group is to renormalize the bottom mass in the scheme de Florian et al. (2016), a calculation in the scheme for both and would be highly desirable. Furthermore, the effect of a finite bottom mass in loops which do not couple to the Higgs, i.e. in the 4-flavor scheme, has seen recent attention in an approximation at leading power in Pietrulewicz and Stahlhofen (2023). As a first rough assessment of such effects, we observe that contributions including -quark PDFs contribute only 0.037 pb to our result at NNLO and 13 TeV. We expect the effect of a switch to the 4-flavor scheme to be of a similar order of magnitude. Still, a complete calculation in this scheme remains of major interest. We defer a dedicated analysis comparing the calculation presented in this letter to an exact calculation in the scheme and/or the 4-flavor scheme to future work. Due to the open questions regarding the scheme choice, we refrain from providing values for top-charm interference, even though the methods used in this work can be readily applied to obtain them. It seems sensible to first address the impact of scheme choice before considering this very small effect, where the usage of the 5-flavor scheme as explained above might no longer be appropriate.